starting from rest a car accelerates at 2.0 m/s^2. How far will it go after 6.0 s


1
Expert's answer
2022-03-15T16:49:41-0400

Explanations & Calculations


  • Since the acceleration is constant apply "\small s=ut+1/2at^2" for its motion.

"\qquad\qquad
\begin{aligned}
\small s&=\small 0+\frac{1}{2}\times2.0\,ms^{-2}\times(6.0\,s)^2\\
&=\small 36.0\,m
\end{aligned}"